The flower garden has the shape of a right triangle. 51 ft of a perennial border forms the hypotenuse of the triangle, and one leg is 21 ft longer than other leg. Find the lengths of the legs. What are the lengths of the legs?
Answer by mananth(16946)

x^2+(x+21)^2=51^2
x^2+x^2+42x+441=2601
2x^2+42x-2160=0
/2
x^2+21x-1080=0
x^2+45x-24x-1080=0
x(x+45)-24(x+45)=0
(x+45)(x-24)=0
x=24 which is positive
one leg = 24 ft
other leg = 45ft